OpenStudy (anonymous):

Find the distance between each pair of parallel lines with the given equations. y=1/3x-3 y=1/3x+2

OpenStudy (anonymous):

Ordered pair (0,-3) belongs to the first equation, Ordered pair (0,2) belongs to the second equation, Formula for distance between 2 points is d=√((x2-x1)^2+(y2-y1)^2) d=√((0-0)^2+(2-(-3) )^2 )=√(5^2 )=5 Distance between 2 parallel lines is 5

OpenStudy (anonymous):

2+3 =5